In a survey conducted by a website, employers were asked if they had ever sent an employee home because they were dressed inappropriately. A total of 2756 employers responded to the survey, with 973 saying that they had sent an employee home for inappropriate attire. In a press release, the website makes the claim that more than one-third of employers have sent an employee home to change clothes.
Do the sample data provide convincing evidence in support of this claim? Test the relevant hypotheses using
α = 0.05.
For purposes of this exercise, assume that it is reasonable to regard the sample as representative of employers in the United States. (Round your test statistic to two decimal places and your P-value to four decimal places.)
z=

P-value=



Answer by Boreal(15235)

This is a one sample proportion test with p hat 973/2756=0.353
z=(phat-p)/SE, where SE is sqrt p*(1-p)/n=sqrt (4/9)/2756; sqrt term=0.009
z=2.20
p-value for one way test=0.014
Reject the null hypothesis that the proportion is the less or the same in favor of the alternative hypothesis.